OpenStudy (anonymous):

use the quadratic formula to solve the equation 5x^2-4x-3=0

OpenStudy (anonymous):

x(5x-4)=3

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

quadratic equation

OpenStudy (anonymous):

can you explain it please?

OpenStudy (amistre64):

you have to know the quadratic formula before you can use it ....

OpenStudy (anonymous):

@DoomDude use quadratic formula not factoring

OpenStudy (amistre64):

the quadratic formula is the aftereffects of completing the square

OpenStudy (anonymous):

can someone explain to me how to use the quadratic formula by using this problem?

OpenStudy (anonymous):

step by step

OpenStudy (amistre64):

\[ax^2+bx+c=0\] \[ax^2+bx=-c\] \[x^2+\frac{b}{a}x=-\frac{c}{a}\] \[x^2+\frac{b}{a}x+(\frac {b}{2a})^2=-\frac ca+(\frac{b}{2a})^2\] \[(x+\frac{b}{2a})^2=-\frac ca+\frac{b^2}{4a^2}\] \[(x+\frac{b}{2a})^2=-\frac{4ac}{4a^2}+\frac{b^2}{4a^2}\] \[x+\frac{b}{2a}=\pm\sqrt{-\frac{4ac}{4a^2}+\frac{b^2}{4a^2}}\] \[x+\frac{b}{2a}=\pm\sqrt{\frac{b^2-4ac}{4a^2}}\] \[x+\frac{b}{2a}=\pm\frac{\sqrt{b^2-4ac}}{2a}\] \[x=-\frac{b}{2a}\pm\frac{\sqrt{b^2-4ac}}{2a}\] \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

bravo, very thorough

OpenStudy (amistre64):

and if you dont want to go thru all the effort of completeing a square; just jump to the end of it and plug in